# Structural and transport properties of LiTFSI/G3 electrolyte with machine-learned molecular dynamics

### v1

@@ Abstract
The lithium bis(trifluoromethylsulfonyl)azanide-triglyme (LiTFSI/G3) electrolyte plays a critical role in the performance of lithium-ion batteries. However, its solvation structure and transport properties at the atomic scale remain incompletely understood. In this study, we develop an efficient and accurate neuroevolution potential (NEP) model by integrating bootstrap and active learning strategies. Using machine-learned NEP-driven molecular dynamics simulations, we explore the structural and diffusion properties of LiTFSI/G3 across a wide range of the solute-to-solvent ratios, systematically analyzing electrolyte density, ion coordination, viscosity, and lithium self-diffusion. The computed densities show excellent agreement with experimental data, and pair correlation analysis reveals significant interactions between lithium ions and surrounding oxygen atoms, which strongly impacts Li+ mobility. Viscosity and diffusion calculations further demonstrate that increasing LiTFSI concentration enhances Li–O interactions, resulting in higher viscosity and reduced lithium diffusion. Additionally, machine learning-based path integral molecular dynamics (PIMD) simulations confirm the negligible impact of quantum effects on Li+ transport. The electrolyte-specific protocol developed in this work provides a systematic framework for constructing high-fidelity machine learning potentials for complex systems.
@@ I Introduction
Lithium-ion batteries (LIBs) are essential for modern energy storage, particularly in electric vehicles and portable electronics, due to their high energy density, long cycle life, and reliability [1]. The electrolyte, a key component, governs ion transport and electrochemical stability, directly influencing battery performance and lifespan [2, 3]. Conventional liquid electrolytes, typically consisting of organic solvents (e.g., carbonates) and lithium salts (e.g., Li+ bis(trifluoromethylsulfonyl)azanide (LiTFSI)), offer high ionic conductivity and electrode compatibility. However, their volatility, flammability, and limited electrochemical stability present significant challenges, necessitating the development of safer and more efficient alternatives.
Molecular dynamics (MD) simulations provide atomic-scale insights into electrolyte structure [4, 5] and transport, revealing key mechanisms of ion solvation, diffusion, and viscosity. However, traditional empirical force fields often fail to accurately capture complex intermolecular interactions, leading to discrepancies in predicting key transport properties. Machine-learned potential (MLP), trained on first-principles data, has emerged as a powerful alternative, achieving near-density functional theory (DFT) accuracy while maintaining computational efficiency. Successfully applied in metals, semiconductors, catalysis [6], and electrolytes [7, 8, 9], MLP leverages large datasets and neural networks to accurately model atomic interactions, providing a robust framework for high-fidelity electrolyte simulations.

### v2

@@ Abstract
The lithium bis(trifluoromethylsulfonyl)azanide-triglyme (LiTFSI/G3) electrolyte plays a critical role in the performance of lithium-ion batteries. However, its solvation structure and transport properties at the atomic scale remain incompletely understood. In this study, we develop an efficient and accurate neuroevolution potential (NEP) model by integrating bootstrap and active learning strategies. Using machine-learned NEP-driven molecular dynamics simulations, we explore the structural and diffusion properties of LiTFSI/G3 across a wide range of the solute-to-solvent ratios, systematically analyzing electrolyte density, ion coordination, viscosity, and lithium self-diffusion. The computed densities show excellent agreement with experimental data, and pair correlation analysis reveals significant interactions between lithium ions and surrounding oxygen atoms, which strongly impacts Li+ mobility. Viscosity and diffusion calculations further demonstrate that increasing LiTFSI concentration enhances Li-O interactions, resulting in higher viscosity and reduced lithium diffusion. Additionally, machine learning-based path integral molecular dynamics (PIMD) simulations confirm the negligible impact of quantum effects on Li+ transport. The electrolyte-specific protocol developed in this work provides a systematic framework for constructing high-fidelity machine learning potentials for complex systems.
@@ I Introduction
Lithium-ion batteries (LIBs) are essential for modern energy storage, particularly in electric vehicles and portable electronics, due to their high energy density, long cycle life [1], and reliability [2, 3]. The electrolyte, a key component, governs ion transport and electrochemical stability [4], directly influencing battery performance and lifespan [5, 6, 7]. Conventional liquid electrolytes, typically consisting of organic solvents (e.g., carbonates) and lithium salts (e.g., Li+ bis(trifluoromethylsulfonyl)azanide (LiTFSI)[8], offer high ionic conductivity and electrode compatibility. However, their volatility, flammability, and limited electrochemical stability present significant challenges, necessitating the development of safer and more efficient alternatives.
Molecular dynamics (MD) simulations provide atomic-scale insights into electrolyte structure [9, 10, 11, 12] and transport [1], revealing key mechanisms of ion solvation, diffusion, and viscosity. However, traditional empirical force fields often fail to accurately capture complex intermolecular interactions, leading to discrepancies in predicting key transport properties. Machine-learned potential (MLP), trained on first-principles data, has emerged as a powerful alternative, achieving near-density functional theory (DFT) accuracy while maintaining computational efficiency. Successfully applied in metals, catalysis [13], and electrolytes [14, 15, 16, 17], MLP leverages large datasets and neural networks to accurately model atomic interactions, providing a robust framework for high-fidelity electrolyte simulations.